Circular No. 6679

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cr IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (V = 20.4, R = 20.4) on CCD images taken on June 7 by S. Chan at the Mount Stromlo 1.27m telescope (+ Macho Camera). SN 1997cr is located near Abell 3744 at α = 21^{h}05^{m}48^{s}.21, δ = −25^{o}13′29′′.8 (equinox 2000.0). A nearby star has position end figures 54^{s}.32, 14′47′′.4. SN 1997cr was also present on similar frames taken on June 1 (same magnitudes), but it is not visible on images taken on 1996 May 22 (limiting mag R = 21). Finding charts are available

at http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

NOVA SCORPII 1997 P. Mason, T. E. Harrison, B. J. McNamara, and J. J. Johnson, New

Mexico State University, report that observations of Liller’s suspected nova (IAUC 6675, 6678) obtained with the ARC 3.5-m telescope confirm that this object is a classical nova. Optical spectroscopy reveals broad emission lines of Hα (FWHM = 3000 km/s), Hβ, the [Ca II] infrared triplet, and strong [Fe II] lines from multiplet 42. The Hβ profile shows evidence for a weak P-Cyg profile.

Z ANDROMEDAE J. A. Mattei, AAVSO, reports that this prototype of symbiotic stars is

undergoing an outburst, as indicated by the following observations submitted to the AAVSO: May 8.46 UT, 10.9 (G. Hanson, II, Cave Creek, AZ); 20.07, 10.6 (G. Poyner, West Midlands, England); 26.36, 10.1 (R. Modic, Richmond Heights, OH); 29.44, 10.0 (Hanson); 31.093, 9.9 (P. Schmeer, Saarbrücken, Germany); June 4.36, 9.7 (Modic); 6.14, 9.8 (J. Bortle, Stormville, NY); 8.36, 9.7 (Hanson); 9.13, 9.5 (Bortle). The last outburst was in 1985 September, when Z And brightened to visual mag 9.1 and was above minimum for ∼ 290 days. Following the 1985 outburst, Z And showed small-amplitude oscillations of ∼ 0.5 mag until the current outburst.

COMET C/1997 J2 (MEUNIER-DUPOUY) Visual m_{1} estimates: May 28.94 UT, 12.0 (A. Pereira, Cabo da Roca,

Portugal, 0.25-m reflector); June 7.93, 12.4 (M. Meyer, Frauenstein, Germany, 0.25-m refl.); 10.13, 12.4 (J. Bortle, Stormville, NY, 0.41-m reflector).

Corrigendum. On IAUC 6659, last line, for 350^{o} read 170^{o}

1997 June 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green